For the reaction 3A(g)+2B(g)⇌C(g) Kc = 92.6 at a temperature of 271 ∘C . Calculate the value of Kp
Answer #1

3A (9) + 2B(9) ec) Here Kc = .92.6 Temperature (T)5 271°C = 2.71+ 273) K = 5544K Now, We have the relation between and ke kp?